# High-Speed Quadrupedal Gaits

### v1

@@ I Introduction
Quadrupedal locomotion has long fascinated researchers in both biology and robotics due to its complexity and adaptability. Among various gaits, galloping stands out as a high-speed strategy characterized by asymmetrical footfall patterns and aerial phases, enabling rapid and agile movement. Understanding the mechanics and energetics of galloping is crucial for developing quadrupedal robots capable of efficient and dynamic locomotion. Galloping is prevalent among mammals such as cheetahs, horses, and dogs. This gait follows a distinct footfall sequence, often incorporating one or more aerial phases per stride. As illustrated by Fig. 1, Hildebrand’s seminal work categorized asymmetrical gaits based on limb phase relationships, distinguishing between transverse gallops, where the hind limb contacts are followed by the ipsilateral forelimb, and rotary gallops, which exhibit a circular limb sequence [1, 2]. These patterns affect speed and stability; for instance, cheetahs employ rotary gallops for sprinting, while horses often use transverse gallops for endurance running [3]. The number of flight phases significantly influences energy expenditure [4, 5]. Galloping with no flight phases is more energy-efficient at lower speeds, while incorporating one or two flight phases enhances efficiency at higher speeds. This adaptability allows animals to optimize gait selection based on environmental and metabolic constraints.
@@
Figure 1: (a) Footfall sequence variations between transverse and rotary galloping gaits, with foot placement color-coded: right-fore foot (blue), left-fore foot (green), right-hind foot (yellow), and left-hind foot (red). (b) Classification of galloping gaits based on suspension phases, including no suspension, gathered suspension, extended suspension, and dual suspension [1]. Animal images sourced from [6].
@@
Translating biological galloping into robotics presents significant challenges. Early efforts focused on walking and trotting, but achieving high-speed, agile locomotion required exploring galloping gaits. The MIT Cheetah robot demonstrated stable galloping using model-based control, reaching speeds of 6 m/s [7, 8]. Another example, the Scout II quadruped, implemented a rotary gallop using passive dynamics and minimal sensory feedback [9], highlighting the role of mechanical design in gait stability. Recent advances have further improved robotic speed and agility. The Black Panther II, developed by Zhejiang University, achieves 10.4 m/s while sprinting 100 meters in under 10 seconds [10]. Similarly, the HOUND robot, developed by KAIST, set a Guinness World Record with a 100-meter sprint time of 19.87 seconds [11]. These developments demonstrate the growing capabilities of quadrupedal robots. However, both robots use trotting, an intermediate-speed gait that prioritizes stability over maximal velocity. In contrast, animals rely on galloping to reach their highest speeds, emphasizing the need for further research into energy-efficient galloping gaits for robotic quadrupeds.
@@
Despite advances in robotic locomotion, replicating efficient galloping remains challenging. Maintaining dynamic stability at high speeds requires precise coordination of limb movements and body dynamics, necessitating robust control strategies. Additionally, robotic galloping often incurs a higher cost of transport (CoT) than biological counterparts, highlighting the need for optimization in both control algorithms and mechanical design. The number of flight phases plays a critical role in energy efficiency. While increased aerial phases enable higher speeds, they also require greater energy input. Optimizing this balance is essential for achieving both speed and efficiency. Furthermore, transitioning between trotting and galloping in unstructured environments introduces additional challenges, as terrain variability affects gait stability.
@@ II-A Footfall Sequence of Galloping Gaits
[…]
@@
  • Foreleg phase lag: $d_{f}\in[-0.5,0.5):=t^{\text{TD}}_{\scriptscriptstyle\text{LF}}-t^{\text{TD}}_{\scriptscriptstyle\text{RF}}$  [^1] Similar modulo operation is used for all three phase lags $d_{f}$, $d_{h}$, and $d_{p}$, for example: $d_{f}:=\left((t^{\text{TD}}_{\scriptscriptstyle\text{LF}}-t^{\text{TD}}_{\scriptscriptstyle\text{RF}}+0.5)\mod 1\right)-0.5$
@@ II-A2 Number of Flight Phases
[…]
GG
:={$:=\{(d_{t},d_{p},d_{f},d_{h})\,|\,d_{p}<d_{t},d_{p}<1-d_{t},\,d_{f}\neq 0,\,d_{h}\neq 0\}$. As indicated in Fig. 3(b), this gait features a single flight phase. During aerial suspension, the legs are gathered under the body, hence a Galloping gait with a Gathered flight phase.
@@
:={$:=\{(d_{t},d_{p},d_{f},d_{h})\,|\,d_{t}<d_{p},1-d_{t}<d_{p},\,d_{f}\neq 0,\,d_{h}\neq 0\}$. As shown in Fig. 3(c), this gait also has a single flight phase. However, during suspension, the legs extend outward from the body, resulting in a Galloping gait with an Extended flight phase.
@@
:={$:=\{(d_{t},d_{p},d_{f},d_{h})\,|\,d_{t}<d_{p}<1-d_{t},\,d_{f}\neq 0,\,d_{h}\neq 0\}$. This gait has two flight phases, as illustrated in Fig. 3(d). Depending on the execution, the torso reaches its highest position either when the legs are gathered or extended, leading to a Galloping gait with 2 flight phases.
@@ II-C Hybrid Dynamics
Galloping in quadrupedal locomotion follows a hybrid dynamical structure, where continuous motion phases are intermittently interrupted by discrete transitions dictated by foot-ground interactions. Specifically, each leg alternates between stance, where the foot maintains ground contact, and swing, when the leg is airborne. These transitions introduce velocity discontinuities, necessitating a mathematical framework that integrates both continuous dynamics and impact-driven state changes, as in [19].
@@ II-D1 Objective Function
[…]
@@ II-D2 Constraints
  • Friction cone: $||{\lambda}_{i,x}+{\lambda}_{i,y}||_{2}-\mu{\lambda}_{i,z}\leq 0$.
@@
Figure 4: CoT as a function of speed for different galloping gaits, comparing rotary (solid lines) and transverse (dashed lines) patterns. G0 is shown in yellow, GE in green, GG in red, and G2 in blue. As speed increases, CoT rises across all gaits, with G0 exhibiting the highest CoT at higher speeds, followed by GG and GE, while G2 maintains the lowest CoT. Minimal differences exist between rotary and transverse variations, indicating similar energy efficiency trends across all gaits.
@@ III-A Cost of Transport
The feasible speed range for each gait differs. G0 is viable between 0.5 m/s and 5.5 m/s, while GG and GE operate from 0.6 m/s to 5.8 m/s. G2 has the broadest range, from 0.5 m/s to 6.6 m/s. A comparison of transverse and rotary galloping reveals nearly identical energy trends, with overlapping CoT curves across all speeds, indicating minimal differences in energy efficiency. This suggests that gait selection between transverse and rotary styles is primarily influenced by stability and biomechanical factors rather than energy consumption. Consequently, we do not differentiate between them in subsequent discussions.
@@ III-C Validating Galloping Gaits in Simulation
To validate our optimal solutions, we conducted simulation tests of the rotary and transverse G2 gaits using the A1 robot in the Gazebo environment, employing a Quadratic Programming (QP)-based locomotion controller developed in our prior work [22]. For each gait, the robot started from rest and gradually accelerated until instability led to a fall at higher speeds. During this process, the CoT was computed stride by stride, as defined in (5), and plotted in Fig. 6(a). As shown in the animations in the accompanying multimedia file, the robot successfully executed galloping gaits with the desired footfall sequences. The CoT trends closely followed those of the optimal solutions in Fig. 4, with rotary and transverse patterns exhibiting nearly identical energy consumption. However, the simulated stride-by-stride CoT was significantly higher than the optimized values. Upon inspection, we found that this discrepancy primarily arose from disturbances in the simulation environment, requiring both the joint feedback controller and foot placement controllers [22] to continuously reject perturbations at high frequencies. This disturbance rejection caused significant joint torque fluctuations, particularly during flight phases, contributing to the observed increase in CoT.
@@ IV Conclusions
The results indicate that gait efficiency is strongly speed-dependent. At low speeds, G0 exhibits the lowest CoT, making it the most energy-efficient option, whereas at higher speeds, G2 becomes the most efficient, maintaining a lower CoT beyond 3 m/s. Minimal differences were observed between transverse and rotary footfall sequences, suggesting that energy efficiency alone does not favor one pattern over the other.
In terms of body motion, roll dynamics are strongly influenced by footfall patterns, particularly in GG, where rotary and transverse gaits exhibit distinct oscillatory behaviors. Large roll variations in GG, GE, and G0 may challenge lateral stability, requiring active control strategies. In contrast, G2 maintains a more stable roll profile, aiding balance in high-speed locomotion. Pitch motion remains consistent across all gaits, indicating it is dictated by overall gait structure rather than footfall sequence. This suggests pitch stabilization can be managed uniformly, simplifying sagittal plane control. Yaw motion is relatively small and negligible in G2, suggesting a minor role in torso dynamics. The reduced yaw oscillations in G2 and GE may enhance heading stability, while GG and G0 exhibit slightly larger yaw deviations, requiring additional control effort. While energy efficiency is speed-dependent, stability considerations favor specific gaits for different tasks. G2 appears optimal for stable, high-speed locomotion, whereas gaits with greater roll and yaw variations may need enhanced stabilization strategies for reliable movement in unstructured environments.

### v2

@@ I Introduction
Quadrupedal locomotion has long fascinated researchers in both biology and robotics due to its complexity and adaptability. Among various gaits, galloping stands out as a high-speed strategy characterized by asymmetrical footfall patterns and aerial phases, enabling rapid and agile movement. Understanding the mechanics and energetics of galloping is crucial for developing quadrupedal robots capable of efficient and dynamic locomotion. Galloping is prevalent among mammals such as cheetahs, horses, and dogs. This gait follows a distinct footfall sequence, often incorporating one or more aerial phases per stride. As illustrated by Fig. 1, Hildebrand’s seminal work categorized asymmetrical gaits based on limb phase relationships, distinguishing between transverse gallops, where the hind limb contacts are followed by the ipsilateral forelimb, and rotary gallops, which exhibit a circular limb sequence [1, 2]. These patterns affect speed and stability; for instance, cheetahs employ rotary gallops for sprinting, while horses often use transverse gallops for endurance running [3]. Additionally, the frequency of the flight phases plays a pivotal role in shaping the overall energy expenditure [4, 5]. Galloping with no flight phases is more energy-efficient at lower speeds, while incorporating one or two flight phases enhances efficiency at higher speeds. This adaptability allows animals to optimize gait selection based on environmental and metabolic constraints.
@@
Figure 1: (a) Footfall sequence variations between transverse and rotary galloping gaits, with foot placement color-coded: right-fore foot (blue), left-fore foot (green), right-hind foot (yellow), and left-hind foot (red). (b) Classification of galloping gaits based on suspension (flight) phases, including no suspension, gathered suspension, extended suspension, and dual suspension [1]. Animal images sourced from [6].
@@
Translating biological galloping into robotics presents significant challenges. Early efforts focused on walking and trotting, but achieving high-speed, agile locomotion required exploring galloping gaits. The MIT Cheetah robots demonstrated stable galloping using model-based control [7, 8]. Another example, the Scout II quadruped, implemented a rotary gallop using passive dynamics and minimal sensory feedback [9], highlighting the role of mechanical design in gait stability. Recent advances have further improved robotic speed and agility. The Black Panther II, developed by Zhejiang University, achieves 10.4 m/s while sprinting 100 meters in under 10 seconds [10]. Similarly, the HOUND robot, developed by KAIST, set a Guinness World Record with a 100-meter sprint time of 19.87 seconds [11]. These developments demonstrate the growing capabilities of quadrupedal robots. However, both robots use trotting, an intermediate-speed gait that prioritizes stability over maximal velocity. In contrast, animals rely on galloping to reach their highest speeds, emphasizing the need for further research into energy-efficient galloping gaits for robotic quadrupeds.
@@
Replicating an efficient galloping gait remains an ongoing challenge. Maintaining dynamic stability at high speeds requires precise coordination of limb movements and body dynamics, necessitating robust control strategies. Additionally, robotic galloping often incurs a higher cost of transport (CoT) than biological counterparts, highlighting the need for optimization in both control algorithms and mechanical design. The number of flight phases plays a critical role in energy efficiency. While increased aerial phases enable higher speeds, they also require greater energy input. Optimizing this balance is essential for achieving both speed and efficiency. Furthermore, transitioning between trotting and galloping in unstructured environments introduces additional challenges, as terrain variability affects gait stability.
@@ II-A Footfall Sequence of Galloping Gaits
The footfall pattern of a gait is uniquely defined by its footfall sequence, represented as $\mathbf{\boldsymbol{t}}=[t^{\text{TD}}_{\scriptscriptstyle\text{LH}},t^{\text{LO}}_{\scriptscriptstyle\text{LH}},t^{\text{TD}}_{\scriptscriptstyle\text{LF}},t^{\text{LO}}_{\scriptscriptstyle\text{LF}},t^{\text{TD}}_{\scriptscriptstyle\text{RF}},t^{\text{LO}}_{\scriptscriptstyle\text{RF}},t^{\text{TD}}_{\scriptscriptstyle\text{RH}},t^{\text{LO}}_{\scriptscriptstyle\text{RH}}]^{T}$. Each timing value $t^{j}_{i}$ (where $i\in\{\text{LH},\text{LF},\text{RF},\text{RH}\}$ and $j\in\{\text{TD},\text{LO}\}$) is normalized by the stride duration $t_{\text{stride}}$ and ranges from 0 to 100%. The initials L, R, F, and H represent left, right, fore, and hind feet, respectively, while TD and LO denote touchdown and lift-off events. Using Hildebrand’s convention [1], footfall sequences are illustrated with black bars for ground contact and white spaces for suspension, with total stride duration spanning the full sequence (e.g., Fig. 2). To quantify different galloping gaits, four phase-based metrics are introduced:
@@
  • Foreleg phase lag: $d_{f}\in[-0.5,0.5):=t^{\text{TD}}_{\scriptscriptstyle\text{LF}}-t^{\text{TD}}_{\scriptscriptstyle\text{RF}}$  [^1] Similar modulo operation is used for all three phase lags $d_{f}$, $d_{h}$, and $d_{p}$, for example: $d_{f}:=\left((t^{\text{TD}}_{\scriptscriptstyle\text{LF}}-t^{\text{TD}}_{\scriptscriptstyle\text{RF}}+0.5)\mod 1\right)-0.5$ .
@@ II-A2 Number of Flight Phases
$\{(d_{t},d_{p},d_{f},d_{h})\,|\,1-d_{t}<d_{p}<d_{t},\,d_{f}\neq 0,\,d_{h}\neq 0\}$. This gallop has two mixed stance phases and no flight phases as shown in Fig. 3(a).
$G$
$\{(d_{t},d_{p},d_{f},d_{h})\,|\,d_{p}<d_{t},d_{p}<1-d_{t},\,d_{f}\neq 0,\,d_{h}\neq 0\}$. As indicated in Fig. 3(b), this gait features a single flight phase. During aerial suspension, the legs are gathered under the body, hence a Galloping gait with a Gathered flight phase.
@@
$\{(d_{t},d_{p},d_{f},d_{h})\,|\,d_{t}<d_{p},1-d_{t}<d_{p},\,d_{f}\neq 0,\,d_{h}\neq 0\}$. As shown in Fig. 3(c), this gait also has a single flight phase. However, during suspension, the legs extend outward from the body, resulting in a Galloping gait with an Extended flight phase.
@@
$\{(d_{t},d_{p},d_{f},d_{h})\,|\,d_{t}<d_{p}<1-d_{t},\,d_{f}\neq 0,\,d_{h}\neq 0\}$. This gait has two flight phases, as illustrated in Fig. 3(d). Depending on the execution, the torso reaches its highest position either when the legs are gathered or extended, leading to a Galloping gait with 2 flight phases.
@@ II-C Hybrid Dynamics
Galloping in quadrupedal locomotion follows a hybrid dynamical structure, where continuous motion phases are intermittently interrupted by discrete transitions dictated by foot-ground interactions. Specifically, each leg alternates between stance, where the foot maintains ground contact, and swing, when the leg is airborne. These transitions introduce velocity discontinuities, necessitating a mathematical framework that integrates both continuous dynamics and impact-driven state changes, as in [18, 19, 17].
@@ II-D1 Objective Function
$\text{CoT}=\frac{\int_{t_{o}}^{t_{f}}|\mathbf{\boldsymbol{u}}\cdot\mathbf{\boldsymbol{\dot{q}}}_{l}|\>dt}{mg\left[\>{x}{\scriptstyle(t_{f})}-{x}{\scriptstyle(t_{o})}\>\right]}$ (5)
@@ II-D2 Constraints
  • Friction cone: ${\lambda}_{i,x}+{\lambda}_{i,y}||_{2}-\mu{\lambda}_{i,z}\leq 0$.
@@
Figure 4: CoT as a function of speed for different galloping gaits, comparing rotary (solid lines) and transverse (dashed lines) patterns. G0 is shown in yellow, GE in green, $G$ in red, and G2 in blue. As speed increases, CoT rises across all gaits, with G0 exhibiting the highest CoT at higher speeds, followed by GG and GE, while G2 maintains the lowest CoT. Minimal differences exist between rotary and transverse variations, indicating similar energy efficiency trends across all gaits.
@@ III-A Cost of Transport
The feasible speed range for each gait differs. G0 is viable between 0.5 m/s and 5.5 m/s, while GG and GE operate from 0.6 m/s to 5.8 m/s. G2 has the broadest range, from 0.5 m/s to 6.6 m/s. A comparison of transverse and rotary galloping reveals nearly identical energy trends, with overlapping CoT curves across all speeds, indicating minimal differences in energy efficiency. This suggests that the choice between transverse and rotary gaits may be driven more by stability, terrain conditions, and biomechanical considerations than by energy consumption. Consequently, we do not differentiate between them in subsequent discussions.
@@ III-C Validating Galloping Gaits in Simulation
To validate our optimal solutions, we conducted simulation tests of the rotary and transverse G2 gaits using the A1 robot in the Gazebo environment, employing a Quadratic Programming (QP)-based locomotion controller developed in our prior work [22]. For each gait, the robot started from rest and gradually accelerated until instability led to a fall at higher speeds. During this process, the CoT was computed stride by stride, as defined in (5), and plotted in Fig. 6(a). As shown in the animations in the accompanying multimedia file, the robot successfully executed galloping gaits with the desired footfall sequences. The CoT trends closely followed those of the optimal solutions in Fig. 4, with rotary and transverse patterns exhibiting nearly identical energy consumption. However, the simulated stride-by-stride CoT was significantly higher than the optimized values. Upon inspection, we found that this discrepancy primarily arose from the behavior of the control system in the simulation environment. Specifically, the joint feedback controller and foot placement controllers [22] responded to frequent perturbations with high-frequency corrections. This led to significant joint torque fluctuations, particularly during flight phases, contributing to the observed increase in CoT.
@@ IV Conclusions
The results indicate that the efficiency of the galloping gaits is highly dependent on speed. At low speeds, G0 exhibits the lowest CoT, making it the most energy-efficient option, whereas at higher speeds, G2 becomes the most efficient, maintaining a lower CoT beyond 3 m/s. This shift may be attributed to the continuous ground contact of G0 at low speeds, which reduces impact forces. In contrast, at higher speeds, galloping gaits like G2, with more frequent flight phases, enable longer strides and leverage ballistic motion, contributing to improved efficiency. Minimal differences were observed between transverse and rotary footfall sequences, suggesting that energy efficiency alone does not favor one pattern over the other. In terms of body motion, roll dynamics are strongly influenced by footfall patterns, particularly in GG, where rotary and transverse gaits exhibit distinct oscillatory behaviors. Large roll variations in GG, GE, and G0 may challenge lateral stability, requiring active control strategies. In contrast, G2 maintains a more stable roll profile, aiding balance in high-speed locomotion. Pitch motion remains consistent across all gaits, indicating it is dictated by overall gait structure rather than footfall sequence. This suggests pitch stabilization can be managed uniformly, simplifying sagittal plane control. Yaw motion is relatively small and negligible in G2, suggesting a minor role in torso dynamics. The reduced yaw oscillations in G2 and GE may enhance heading stability, while GG and G0 exhibit slightly larger yaw deviations, requiring additional control effort. While energy efficiency is speed-dependent, stability considerations favor specific gaits for different tasks. G2 appears optimal for stable, high-speed locomotion, whereas gaits with greater roll and yaw variations may need enhanced stabilization strategies for reliable movement in unstructured environments.